OpenStudy (anonymous):

If flipping a coin is equally likely to result in heads or tails, the odds of flipping a coin six times and getting only one head are 3 out of 16. A. True B. False

pooja195 (pooja195):

If you flip a coin 6 times, there are 2 possibilities for each coin, so the total number of possibilities is 2^6 = 64, and we are assuming they are equally as likely. There are six ways that you can get just one head. The first coin could come up a head with the rest tails, or the second one can be the only head, and so on for each of the 6 coins. The probability of only one coin being a head is therefore 6/64 = 3/32, which is not 3/16, so the answer is False.